OpenStudy (anonymous):

the line passing through (9,y) and (3,0) is parallel to the joining (3,4) and (-1,2). find y? i got 3 is this correct?

OpenStudy (mathstudent55):

Slope of line passing through (3, 4) and (-1, 2) m = (2 - 4)/(-1 - 3) = -2/-4 = 1/2 Slope of line passing through (9, y) and (3, 0) m = (y - 0)/(9 - 3) = 1/2 y/6 = 1/2 y = 3 You are correct.

OpenStudy (anonymous):

thanks :)

OpenStudy (mathstudent55):

wlcm